# An Embedding is Worth a Thousand Noisy Labels

### v1

@@ 4.7 Dimensionality reduction
To showcase the effectiveness of the proposed Filtered LDA (FLDA) approach, we evaluate WANN’s classification accuracy on the linearly projected test set, utilizing the complete projected training set as WANN’s support set. Four different projections are employed for this analysis. First, as a baseline, we report classification performances without dimensionality reduction, labeled as “None”. Subsequently, we utilize Principal Component Analysis (PCA) with a fixed number of components, specifically $200$, explaining an average of $77.19\%$ of the total variance across each dataset. Additionally, we compare this with plain Linear Discriminant Analysis (LDA) fed with the entire noisy dataset. Finally, we present the effectiveness of our filtering approach before applying LDA, named as FLDA. To this extent, we recall that LDA projects the dataset along $C-1$ axes, where $C$ is the number of classes. Consistently with our previous experimental settings, the accuracy is reported on datasets with 60% symmetric noise, 30% asymmetric noise, and, 40% instance-dependent label noise. The experiments cover four different datasets: CIFAR-10, CIFAR-100 [25], MNIST [27], and Fashion-MNIST [51]. Note that all these datasets consist of 10 classes, except for CIFAR-100, which includes 100 classes.
Figure 6: Classification accuracy of WANN under three dimensionality reduction strategies, evaluated across four datasets and three different noise settings each. None reflects the performance using the original image embeddings without dimensionality reduction. PCA and LDA show the classification performances on the linearly projected feature spaces, utilizing the whole noisy training set. Finally, FLDA involves the proposed filtering approach. FLDA is the significantly best dimensionality reduction method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).
Results   The results are visually reported in Figure 6. As expected, PCA with $200$ components yields marginal improvement compared to the baseline, suggesting the need for a more nuanced linear projection. Surprisingly, LDA trained on the entire noisy label set already exhibits improved performance on some datasets relative to the baseline. This improvement could be attributed to the significantly reduced dimensionality, minimizing overall sparsity. Therefore, large neighborhoods may compensate for imperfect projections. However, FLDA is significantly better than all other techniques evaluated (Wilcoxon signed-rank test, $p{<}0.05$). The reduction in dimensionality not only enhances classification performance but also contributes to faster computational speed and potential storage efficiency. Notably, using DINOv2 ViT-L as feature extractor, we have an initial embedding size of $1024$. Thus, we significantly improved the classification performance with 100×100$\times$100 × and 10×10$\times$10 × smaller embeddings.
@@ 4.8 Explainability benefits
While deep learning methods are often viewed as black boxes with limited explainability, some traditional machine learning approaches inherently allow for a certain level of explainability. Notably, $k$-NN-based methods provide explainability through the representative nature of the neighborhood. Exploring the neighborhood for each prediction provides a deeper understanding of the factors contributing to correct or erroneous classifications. As illustrated in Figure 2, we present four test samples along with their top $3$ closest training examples from STL-10 [8], a subset of ImageNet [13]. In Figure 2(a), we highlight two test images with clearly incorrect reference labels, supported by clean training samples. In contrast, Figure 2(b) shows two ambiguous test labels, including multiple known objects within a single image, such as a bird with a ship in the background or a dog playing with a cat. The pitfalls of ImageNet labels were already illustrated by [4]. However, this provides a simple yet effective way to manually inspect potential wrong labels within our datasets. Indeed, the explainability of $k$-NN-based approaches becomes particularly valuable in the context of active label correction [36, 2], where human intervention is employed to re-annotate noisy samples. Beyond identifying noisy samples, these methods offer the potential to provide a reliable soft-label candidate derived from the majority of the neighborhood. Additionally, visualizing the neighbors can provide more evidence, especially for challenging or overlapping classes.
@@ 5 Discussion and conclusions
Limitations   The effectiveness of $k$-NN relies on a representative feature representation, posing challenges for domains not represented during pre-training. However, this limitation can be mitigated by using alternative feature extractors. This is now possible thanks to the growing availability of such large open-source models. Moreover, while the proposed FLDA significantly enhances classification performance, it requires a sufficient sample size to generate reliable projection vectors. Finally, to mitigate the computational demands of $k$-NN’s exhaustive search, established Approximate-NN methods [32, 17, 21] provide an effective solution, achieving comparable classification performance with substantially lower computational complexity. Indeed, we could expect a negligible influence on classification performance, as distance metrics inherently only approximate the nearest sample.
Side benefits and generalization   The efficiency of our classification method, further optimized with dimensionality reduction, enables its deployment with limited resources. Moreover, the reliable feature space of current feature extractors becomes particularly suitable for $k$-NN-based approaches, possibly generalizing to different domains. For instance, we observed outstanding classification and robustness performance in the medical domain, where interpretability, data scarcity, and data imbalance are still open challenges. Lastly, the compliance with recent regulations, such as the right to be forgotten [30], is facilitated by the proposed database approach. The flexibility to remove or add data points without extensive re-training or fine-tuning, aligns with privacy guidelines, providing a cost-effective solution.
Conclusion   We introduce WANN, a Weighted Adaptive Nearest Neighbor approach, relying on a reliability score quantifying the correctness of the training labels. Our experiments demonstrate its higher robustness with respect to label noise, particularly in scenarios with limited labeled data, surpassing traditional models relying on robust loss functions. Additionally, the incorporation of dimensionality reduction techniques enhances its performance under noisy labels, facilitated by the proposed weighted scoring.
Data Availability Statement
All experiments are conducted on publicly available datasets; see the references cited.

### v3

@@ 4.7 Dimensionality reduction
To showcase the effectiveness of the proposed Filtered LDA (FLDA) approach, we evaluate WANN’s classification accuracy on the linearly projected test set, utilizing the complete projected training set as WANN’s support set. Four different projections are employed for this analysis. First, as a baseline, we report classification performances without dimensionality reduction, labeled as “None”. Subsequently, we utilize Principal Component Analysis (PCA) with a fixed number of components, specifically $200$, explaining an average of $77.19\%$ of the total variance across each dataset. Additionally, we compare this with plain Linear Discriminant Analysis (LDA) fed with the entire noisy dataset. Finally, we present the effectiveness of our filtering approach before applying LDA, named as FLDA. To this extent, we recall that LDA projects the dataset along $C-1$ axes, where $C$ is the number of classes. Consistently with our previous experimental settings, the accuracy is reported on datasets with 60% symmetric noise, 30% asymmetric noise, and, 40% instance-dependent label noise. The experiments cover three different datasets: CIFAR-10, CIFAR-100, and MNIST (Lecun et al., 1998). Following Ye et al. (2023), for asymmetric noise on CIFAR-100 we group the 100 classes into 20 super-classes and flip each class within its super-class in a circular fashion, while on MNIST we use the mapping $7\rightarrow 1$, $2\rightarrow 7$, $5\leftrightarrow 6$, $3\rightarrow 8$.
Figure 6: Classification accuracy of WANN under three dimensionality reduction strategies, evaluated across three datasets and three different noise settings each. None reflects the performance using the original image embeddings without dimensionality reduction. PCA and LDA show the classification performances on the linearly projected feature spaces, utilizing the whole noisy training set. Finally, FLDA involves the proposed filtering approach. FLDA is the significantly best dimensionality reduction method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).
Results   The results are visually reported in Figure 6. On the clean dataset, LDA and FLDA exhibit among the highest gains, substantially reducing the dimensionality and thereby reducing the overall sparsity. Here, LDA outperforms its filtered counterpart because of the absence of noisy labels. In fact, these performance gains vanish in the presence of label noise, confirming that FLDA is significantly better than all other dimensionality-reduction methods (Wilcoxon signed-rank test, $p{<}0.05$) under varying noise conditions. The reduction in dimensionality not only enhances classification performance but also contributes to faster computational speed and potential storage efficiency.
Notably, using DINOv2 ViT-L as feature extractor, we have an initial embedding size of $1024$. Thus, we significantly improved the classification performance with 100×100$\times$100 × (CIFAR-10, MNIST) and 10×10$\times$10 × (CIFAR-100) smaller embeddings.
@@ 4.8 Explainability benefits
While deep learning methods are often viewed as black boxes with limited explainability, some traditional machine learning approaches inherently allow for a certain level of explainability. Notably, $k$-NN-based methods provide explainability through the representative nature of the neighborhood. Exploring the neighborhood for each prediction provides a deeper understanding of the factors contributing to correct or erroneous classifications. As illustrated in Figure 2, we present four test samples along with their top $3$ closest training examples from STL-10 (Coates et al., 2011), a subset of ImageNet (Deng et al., 2009). In Figure 2(a), we highlight two test images with clearly incorrect reference labels, supported by clean training samples. In contrast, Figure 2(b) shows two ambiguous test labels, including multiple known objects within a single image, such as a bird with a ship in the background or a dog playing with a cat. The pitfalls of ImageNet labels were already illustrated by Beyer et al. (2020). However, this provides a simple yet effective way to manually inspect potentially wrong labels within our datasets. Indeed, the explainability of $k$-NN-based approaches becomes particularly valuable in the context of active label correction (Rebbapragada et al., 2012; Bernhardt et al., 2022), where human intervention is employed to re-annotate noisy samples.
@@ 5 Discussion and conclusions
Limitations   The effectiveness of $k$-NN relies on a representative feature representation, posing challenges for domains not represented during pre-training. However, this limitation can be mitigated by using alternative feature extractors. This is now possible thanks to the growing availability of such large open-source models. Moreover, while the proposed FLDA significantly enhances classification performance, it requires a sufficient sample size to generate reliable projection vectors. Finally, to mitigate the computational demands of $k$-NN’s exhaustive search, established Approximate-NN methods (Malkov & Yashunin, 2020; Guo et al., 2020; Johnson et al., 2021) provide an effective solution, achieving comparable classification performance with substantially lower computational complexity. Indeed, we could expect a negligible influence on classification performance, as distance metrics inherently only approximate the nearest sample.
Side benefits and generalization   The efficiency of our classification method, further optimized with dimensionality reduction, enables its deployment with limited resources. Moreover, the reliable feature space of current feature extractors becomes particularly suitable for $k$-NN-based approaches, possibly generalizing to different domains. For instance, we observed outstanding classification performance and robustness in the medical domain, where interpretability, data scarcity, and data imbalance are still open challenges. Furthermore, the compliance with data regulations, such as the right to be forgotten (Mantelero, 2013), is facilitated by the proposed database approach. The flexibility to remove or add data points without extensive re-training or fine-tuning, aligns with privacy guidelines, providing a cost-effective solution. Lastly, our method can be readily applied to any domain, tabular or otherwise, as long as we can obtain a meaningful semantic space. While our current work focuses on image embeddings derived from large vision foundation models, the same concept extends to language, audio, or multimodal embeddings, assuming they capture the semantic relationships necessary for defining reliable neighborhoods.
Conclusion   We introduce WANN, a Weighted Adaptive Nearest Neighbor approach, based on a reliability score quantifying the correctness of the training labels. Our experiments demonstrate its higher robustness with respect to label noise, particularly in scenarios with limited labeled data, surpassing traditional models relying on robust loss functions. Additionally, the incorporation of dimensionality reduction techniques enhances its performance under noisy labels, facilitated by the proposed weighted scoring.
@@ Acknowledgments
Appendix A Generalizability across backbones
This section illustrates two additional experiments using a ViT-L backbone pretrained on ImageNet in a supervised fashion. We compared our method with the same subset of loss functions utilized in Section 4.4 (CE, ELR, NCE-AGCE, ANL-FL, ANL-CE-ER). On CIFAR-N (c.f. Table A.1), while WANN shows slightly lower performance than the best-performing robust loss function on this particular backbone, our FLDA approach offers a substantial performance boost, achieving results comparable to or even surpassing those of the robust loss functions.
Additionally, we conducted an experiment analogous to Section 4.3 using CIFAR-10 with subsets of $50$ and $100$ samples per class (c.f. Table A.2). Notably, when using only $50$ samples per class, the adaptive methods substantially outperform the robust loss functions. These findings demonstrate that the advantages of our approach are not merely due to the DINOv2 backbone’s inherent strengths but rather underscore the general applicability and robustness of our method.
CIFAR-10N CIFAR-100N
Noisy split Clean Aggr. R1 R2 R3 Worst Clean Noisy
NR ($\approx$) - 9.01% 17.23% 18.12% 17.64% 40.21% - 40.20%
CE 96.56$\pm$0.06 95.19$\pm$0.09 94.96$\pm$0.27 95.01$\pm$0.16 95.03$\pm$0.15 91.35$\pm$0.16 84.66$\pm$0.30 74.48$\pm$0.10
ELR 96.37$\pm$0.09 95.94$\pm$0.07 95.74$\pm$0.14 95.85$\pm$0.07 95.71$\pm$0.38 94.25$\pm$0.35 84.03$\pm$0.06 76.50$\pm$0.28
NCE-AGCE 96.50$\pm$0.06 96.10$\pm$0.07 95.96$\pm$0.06 95.98$\pm$0.10 95.93$\pm$0.09 93.79$\pm$0.25 84.62$\pm$0.17 77.92$\pm$0.09
ANL-FL 96.09$\pm$0.19 95.90$\pm$0.10 95.60$\pm$0.23 95.78$\pm$0.10 95.45$\pm$0.27 94.41$\pm$0.12 79.41$\pm$0.33 71.24$\pm$0.32
ANL-CE-ER 96.18$\pm$0.14 95.78$\pm$0.26 95.52$\pm$0.35 95.63$\pm$0.32 95.72$\pm$0.16 94.79$\pm$0.15 79.57$\pm$0.30 73.07$\pm$0.26
ANN 94.97 94.50 94.25 94.05 94.19 90.53 78.31 70.29
WANN 94.64 94.36 94.15 94.17 94.25 92.22 77.61 70.95
WANN+FLDA 95.99 95.81 95.77 95.86 95.90 92.94 81.93 75.15
Table A.1: Accuracy on CIFAR-N datasets utilizing a ViT Large pre-trained on ImageNet. The results are averaged over five seed runs.
CIFAR-10 (#$\#50$) CIFAR-10 (#$\#100$)
NR - 60% 30% 40% - 60% 30% 40%
CE 95.23$\pm$0.67 74.46$\pm$2.13 88.07$\pm$1.45 81.76$\pm$1.96 95.95$\pm$0.46 83.87$\pm$1.00 90.92$\pm$0.77 86.89$\pm$2.02
ELR 95.00$\pm$0.34 82.40$\pm$4.65 93.63$\pm$0.63 89.98$\pm$4.15 96.45$\pm$0.49 94.02$\pm$1.90 94.14$\pm$2.14 94.32$\pm$1.81
NCE-AGCE 95.49$\pm$0.97 85.94$\pm$2.62 93.22$\pm$1.13 90.82$\pm$3.19 96.93$\pm$0.35 94.71$\pm$0.96 95.66$\pm$0.57 95.40$\pm$0.84
ANL-FL 94.39$\pm$1.08 78.00$\pm$2.31 88.02$\pm$2.77 82.80$\pm$2.15 96.26$\pm$0.70 90.54$\pm$1.40 92.10$\pm$0.65 88.79$\pm$2.24
ANL-CE-ER 94.82$\pm$0.89 78.18$\pm$3.90 90.05$\pm$0.71 83.13$\pm$4.20 95.99$\pm$0.76 90.54$\pm$0.97 92.97$\pm$1.06 89.20$\pm$1.89
ANN 97.57$\pm$0.14 94.00$\pm$0.62 94.48$\pm$0.82 93.75$\pm$1.79 88.70$\pm$1.32 96.74$\pm$0.37 92.37$\pm$1.19 94.47$\pm$1.66
WANN 97.58$\pm$0.15 95.47$\pm$0.68 95.32$\pm$0.74 95.22$\pm$1.19 88.69$\pm$1.31 97.10$\pm$0.36 92.67$\pm$1.26 94.85$\pm$1.55
Table A.2: Accuracy on limited noisy data ($50$ and $100$ samples per class) utilizing a ViT Large pre-trained on ImageNet. The results are averaged over five seed runs.
Appendix B Robustness of reliability score
B.1 Signal for linear classifiers
This experiment aims to evaluate the utility and robustness of our reliability score. The hypothesis is the following: if our reliability score is a robust indicator of label noise, a linear layer should be able to utilize this information just as effectively as $k$-NN. To test this hypothesis, we precomputed the reliability score for each training sample and concatenated it with the respective feature embedding. For test samples, we utilized an average reliability score from their training neighborhood, based on the intuition that the spatial distribution of these scores is both informative and reliable.
As illustrated in Table B.1, for CIFAR-10 and CIFAR-100, this straightforward approach of augmenting the feature embeddings with the reliability score slightly improves performance while further reducing the variability (lower standard deviation across five seed runs). These results hold promise that the reliability score provides a valuable signal that a linear classifier can utilize as effectively as the $k$-NN approach.
CIFAR-10 CIFAR-100
NR - 60% 30% 40% - 60% 30% 40%
CE 99.40$\pm$0.03 98.00$\pm$0.13 96.47$\pm$0.19 95.66$\pm$0.72 78.82$\pm$0.32 60.17$\pm$2.02 69.09$\pm$1.53 54.57$\pm$4.89
$\eta$-CE 99.36$\pm$0.02 98.10$\pm$0.13 96.55$\pm$0.31 95.94$\pm$0.51 80.22$\pm$1.71 63.48$\pm$2.01 70.11$\pm$1.18 60.04$\pm$3.29
Table B.1: Accuracy on CIFAR-10 and CIFAR-100 under three label noise settings. We compare the standard performance using the raw embeddings (CE) and the ones concatenated with our reliability score ($\eta$-CE). The results are averaged over five seed runs.
B.2 Distribution of the reliability scores
Utilizing a toy dataset consisting of $1,000$ samples and $40\%$ symmetric noise, we illustrate in Figure B.1: (top-left) the real class distribution, (top-right) the noisy class distribution, (bottom-left) the reliability scores assigned to each point, and (bottom-right) the distribution of reliability scores for clean vs. noisy samples. From the bottom-right plot, it is possible to observe that clean samples (blue) tend to have higher reliability scores, whereas noisy samples (red) cluster at lower or intermediate values. Points near decision boundaries or mislabeled neighbors also receive moderate or low reliability scores. This observation supports our intuition that WANN utilizes local consistency in the embedding space to effectively distinguish clean from noisy samples.
Figure B.1: Reliability score distribution on a toy dataset with $40\%$ symmetric noise.